OpenStudy (anonymous):

In an isolated system, two roller skaters, each with a mass of 100 kg, collide. Skater 1 was initially at rest, while Skater 2 was moving at 6 m/s. They move off together. What is their speed?

OpenStudy (jamesj):

You need to to use conservation of momentum.

OpenStudy (jamesj):

Let p1 be the momentum of the first skater before the collision on p1' afterwards. Likewise p2 and p2' for skater 2. Then by conservation of momentum \[ p_1 + p_2 = p_1' + p_2' \] Now, use that the definition of momentum to find the joint velocity afterwards.

OpenStudy (anonymous):

so 6m/s?

OpenStudy (jamesj):

what's the definition of momentum?

OpenStudy (anonymous):

mass x velocity

OpenStudy (jamesj):

Yes. So p1 = ? and p2 = ?

OpenStudy (anonymous):

0 and 6ms?

OpenStudy (jamesj):

\[ p_1 = 0 \ kg.m/s \] yes. But \[ p_2 = m_2.v_2 = ... what? \]

OpenStudy (anonymous):

10m/s?

OpenStudy (jamesj):

what's m2 = ? what's v2 = ?

OpenStudy (anonymous):

3 m/s?

OpenStudy (jamesj):

\[ v_2 = 6 m/s \] \[ m_2 = 100 kg \] Hence \[ p_2 = m_2.v_2 = (100 kg)(6 m/s) = 6000 kg.m/s \] Therefore the total momentum before the collision is \[ p_1 + p_2 = 0 + 6000 = 6000 \] Now what about afterwards? What do you know about v1' and v2' ?

OpenStudy (anonymous):

is it 6 m/s...that's all i need to know..

OpenStudy (jamesj):

No. That's wrong. After the collision, they both move off together so their velocities are the same. Call it v: \[ v = v_1' = v_2' \] Hence the total momentum after the collision is \[ p_1' + p_2' = m_1v_1' + m_2v_2' = (m_1 + m_2)v \]

OpenStudy (anonymous):

0!

OpenStudy (jamesj):

Now, set that equal to the total momentum before the collision and you can solve for v.

OpenStudy (anonymous):

0 m/s!

OpenStudy (jamesj):

No. That can't possibly be right. If they are both stationary, v = 0, then their total momentum is zero. But we know it must be equal to 6000.

OpenStudy (anonymous):

so 10??

OpenStudy (jamesj):

Work the algebra here. Set that last expression for the total momentum equal to the total momentum before the collision.

OpenStudy (anonymous):

it has to be one of them, its a multiple choice question.

OpenStudy (anonymous):

can you please just tell me the answer i'll ask my teacher tomorrow

OpenStudy (jamesj):

You're just guessing the answers. I'm trying to show you how to find it. By Conservation of Momentum, momentum before and after the collision are equal: \[ p_1 + p_2 = p_1' + p_2' \] We know that \[ p_1 + p_2 = 6,000 \ kg.m/s \] We also know that \[ p_1' + p_2' = (m_1 + m_2)v \] where v is their shared velocity after the collision. Hence we have that \[ (m_1 + m_2)v = 6000 \] Now solve for v.

OpenStudy (jamesj):

**correction: p1 + p2 = 600, because m2.v2 = (100)(6) = 600. Hence you have that \[ (m_1 + m_2)v = 600 \]

OpenStudy (anonymous):

10 and 6

OpenStudy (jamesj):

Now, what's the value of \[ m_1 + m_2 \]?

OpenStudy (anonymous):

well one is 6...so the other has to be 10!

OpenStudy (jamesj):

they both weight 100 kg. Hence \[ m_1 + m_2 = ... what? \] and hence what is the solution to \[ (m_1 + m_2)v = 600 \ kg.m/s \]

OpenStudy (anonymous):

if they weigh 100, it must be 6

OpenStudy (jamesj):

\[ m_1 = m_2 = 100 \ kg \] hence \[ m_1 + m_2 = 200 \ kg \] therefore \[ (200 \ kg) v = 600 \ kg.m/s \] hence \[ v = ...what? \]

OpenStudy (anonymous):

3!!!!!!!

OpenStudy (jamesj):

v = 3 m/s, yes. Makes intuitive sense as well, as that's half the original velocity.

OpenStudy (anonymous):

god i'm stupid. thank you!

OpenStudy (jamesj):

Now, do yourself a favor, and write this solution out again on a blank piece of paper. When you can do that without looking at this or anything else, then you understand the problem.

OpenStudy (anonymous):

just did. thank you!